# Diffusion approximation of the stationary distribution of a two-level single server queue

### v6

@@ 4.2 Preliminary lemmas
The first equation of (4.45) is immediate from (4.2) and Lemma 4.6 together with (4.43) and (4.44), and the second equation follows from the first one by Lemma 4.1. As for (4.46), we only prove its first equation because its second one is similarly proved. By the Schwartz’s inequality,
$\big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{2}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\big{)}^{2}=\big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{3/2}(\widehat{R}^{(r)}_{s})^{1/2}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\big{)}^{2}$
$\leq\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{3}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})].$ (4.47)
From (4.5) for $k=3$ and Lemma 4.4, we have
$\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{3}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\leq\mathbb{E}[(T^{(r)}_{s})^{3}],$
which is uniformly bounded for $r\in(0,1]$, and therefore (4.3) and (4.44) prove (4.46). ∎
(2nd Step)   In this step, we first expand the exponential terms in $D^{(r)}(\theta_{1},\theta_{2})$ of (4.2), using Lemmas 4.7 and 4.8 in the step 1, then
$r^{-1}D^{(r)}(\theta_{1},\theta_{2})$
$=\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}-e^{\theta_{1}\ell_{0}}+o(r))\mathbb{E}^{(r)}_{e}[(r^{-1}+\mu\theta_{2}\widehat{R}^{(r)}_{s}-2^{-1}r\mu^{3}\sigma_{s}^{2}\theta_{2}^{2}(\widehat{R}^{(r)}_{s})^{2})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(e^{\theta_{1}\ell_{0}}+o(r))\mathbb{E}^{(r)}_{e}[(\mu(\theta_{2}-\theta_{1})\widehat{R}^{(r)}_{s}-2^{-1}r\mu^{3}\sigma_{s}^{2}(\theta_{2}^{2}-\theta_{1}^{2})(\widehat{R}^{(r)}_{s})^{2})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+r(e^{\theta_{1}\ell_{0}}\theta_{1}-e^{\theta_{2}\ell_{0}}\theta_{2})+o(r))$
×$\hskip 51.6665pt\times\mathbb{E}^{(r)}_{s}[(r^{-1}-\lambda_{1}\theta_{1}\widehat{R}^{(r)}_{e}-2^{-1}r\lambda_{1}^{3}\sigma_{e,1}^{2}\theta_{1}^{2}(\widehat{R}^{(r)}_{e})^{2})1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}(1+\theta_{2}r)+o(r))\mathbb{E}^{(r)}_{s}[((\lambda_{2}\theta_{2}-\lambda_{1}\theta_{1})\widehat{R}^{(r)}_{e}$
$\hskip 51.6665pt+2^{-1}r(\lambda_{2}^{3}\sigma_{e,2}^{2}r\theta_{2}^{2}-\lambda_{1}^{3}\sigma_{e,1}^{2}r\theta_{1}^{2})(\widehat{R}^{(r)}_{e})^{2})1(L^{(r)}(0)=\ell^{(r)}_{0})]+o(r).$ (4.48)
We next apply Lemma 4.9 to this formula, then
$r^{-1}D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}-e^{\theta_{1}\ell_{0}}+o(r))\mathbb{E}^{(r)}_{e}[(r^{-1}+\mu\theta_{2}\widehat{R}^{(r)}_{s})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(e^{\theta_{1}\ell_{0}}+o(r))\mathbb{E}^{(r)}_{e}[\mu(\theta_{2}-\theta_{1})\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+r(e^{\theta_{1}\ell_{0}}\theta_{1}-e^{\theta_{2}\ell_{0}}\theta_{2})+o(r))\mathbb{E}^{(r)}_{s}[(r^{-1}-\lambda_{1}\theta_{1}\widehat{R}^{(r)}_{e})1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}(1+\theta_{2}r)+o(r))\mathbb{E}^{(r)}_{s}[(\lambda_{2}\theta_{2}-\lambda_{1}\theta_{1})\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]+O(r)$
$=\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}\theta_{2}-e^{\theta_{1}\ell_{0}}\theta_{1})(\mu\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}])$
$\quad+\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}\lambda_{2}\theta_{2}-e^{\theta_{1}\ell_{0}}\lambda_{1}\theta_{1})\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]+o(r).$ (4.49)
Recall that
$\Delta^{(r)}_{i}=\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\lambda^{(r)}_{i}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\hskip 43.05542pt-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}],\qquad i=1,2,$
then (4.3) can be written as
$D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}r\big{(}e^{\theta_{2}\ell_{0}}\theta_{2}\Delta^{(r)}_{2}-e^{\theta_{1}\ell_{0}}\theta_{1}\Delta^{(r)}_{1}\big{)}+o(r^{2}).$ (4.50)
Substituting this into BAR (4.3) and dividing by $r$, we have
$\frac{1}{2}\sum_{i=1,2}r\mu c_{i}\sigma_{i}^{2}\big{(}\beta_{i}-\theta_{i}\big{)}\theta_{i}\psi^{(r)}_{i}(\theta_{i})$
$\quad=c_{1}\mu\theta_{1}\mathbb{P}[L^{(r)}=0]+\alpha^{(r)}_{e}\big{(}e^{\theta_{2}\ell_{0}}\theta_{2}\Delta^{(r)}_{2}-e^{\theta_{1}\ell_{0}}\theta_{1}\Delta^{(r)}_{1}\big{)}+o(r).$ (4.51)
Substituting $\theta_{2}=0$ or $\theta_{1}=0$, this asymptotic BAR is split into two equations:
$\frac{1}{2}r\mu c_{1}\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})=c_{1}\mu\mathbb{P}[L^{(r)}=0]-\alpha^{(r)}_{e}e^{\theta_{1}\ell_{0}}\Delta^{(r)}_{1}+o(r),$ (4.52)
$\frac{1}{2}r\mu c_{2}\sigma_{2}^{2}\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})=\alpha^{(r)}_{e}e^{\theta_{2}\ell_{0}}\Delta^{(r)}_{2}+o(r).$ (4.53)
(3rd Step)   We compute $\psi^{(r)}_{i}(\theta_{i})$ for $i=1,2$ from (4.52) and (4.53) separately. First consider (4.53). Note that (4.34) yields
$\alpha^{(r)}_{e}\Delta^{(r)}_{2}=r\mu b_{2}\varphi^{(r)}_{2}(0)+o(r).$ (4.54)
Hence, from (4.53), we have
$\frac{1}{2}r\mu c_{2}\sigma_{2}^{2}\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})$ $=\frac{1}{2}e^{\theta_{2}\ell_{0}}r\mu c_{2}\sigma_{2}^{2}\beta_{2}\varphi^{(r)}_{2}(0)+o(r),$
which is simply written as
$\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})$ $=e^{\theta_{2}\ell_{0}}\beta_{2}\varphi^{(r)}_{2}(0)+o(1).$ (4.55)
We next consider (4.52). From (4.33) of Lemma 4.6,
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=-r\mu b_{1}\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]+c_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]+o(r).$ (4.56)
Hence, from (4.52), we have
$r\mu c_{1}\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})$
$\qquad=2c_{1}\mu(1-e^{\theta_{1}\ell_{0}})\mathbb{P}[L^{(r)}=0)]+2r\mu b_{1}e^{\theta_{1}\ell_{0}}\varphi^{(r)}_{1}(0)+o(r).$ (4.57)
This yields
@@
$\mathbb{P}[L^{(r)}=0)]=r\frac{c_{1}\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})-2b_{1}e^{\theta_{1}\ell_{0}}\varphi^{(r)}_{1}(0)}{2c_{1}(1-e^{\theta_{1}\ell_{0}})}+o(r).$ (4.58)
We separately consider (4.58) for $b_{1}=0$ and $b_{1}\not=0$.
(Case of $b_{1}=0$)   In this case, $\beta_{1}=b_{1}=0$, and therefore (4.58) yields
$\mathbb{P}[L^{(r)}=0)=r\lim_{\theta_{1}\to 0}\frac{r\sigma_{1}^{2}\theta_{1}\psi^{(r)}_{1}(\theta_{1})}{2(e^{\theta_{1}\ell_{0}}-1)}+o(r)=\frac{r\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)+o(r).$ (4.59)
Hence, for $b_{1}=0$, it follows from (4.56) and (4.59) that
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\frac{r\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)+o(r),$ (4.60)
$\psi^{(r)}_{1}(\theta_{1})=\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}}\varphi^{(r)}_{1}(0)+o(r).$ (4.61)
@@
(Case of $b_{1}\not=0$)   By substituting $\theta_{1}=\beta_{1}$ into (4.58), we have
$\mathbb{P}[L^{(r)}=0)]=\frac{rb_{1}e^{\beta_{1}\ell_{0}}}{c_{1}(e^{\beta_{1}\ell_{0}}-1)}\varphi^{(r)}_{1}(0)+o(r)$ (4.62)
and therefore, it follows from (4.56), (4.3) and (4.62) that
@@
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\frac{r\mu b_{1}}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(r)}_{1}(0)+o(r),$ (4.63)
@@
$\psi^{(r)}_{1}(\theta_{1})=\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\varphi^{(r)}_{1}(0)+o(r).$ (4.64)
(4th Step)   We now complete the proof of Theorem 3.1. We first prove (i). Note that the bounded away assumption can be written as
$\liminf_{r\downarrow 0}\varphi^{(r)}_{i}(0)>0,\qquad i=1,2.$ (4.65)
Under this condition, for $\theta_{1}\in\mathbb{R}_{+}$ and $\theta_{2}\leq 0$,
@@
is equivalent to
$\lim_{r\downarrow 0}|(\psi^{(r)}_{i}(\theta_{1})-\varphi^{(r)}_{i}(\theta_{1}))|=0,$ (4.67)
since $\varphi^{(r)}_{i|cd}(\theta_{i})=\varphi^{(r)}_{i}(\theta_{i})/\varphi^{(r)}_{i}(0)$. Here, (4.67) holds because (4.13) implies that
Hence, (4.66) holds. Applying it to (4.61) and (4.64), we obtain
@@
Since $\widetilde{\varphi}_{i|cd}$ is the moment generating function of $\widetilde{\nu}_{i|cd}$, (4.68) and (4.69) are equivalent to that $\nu^{(r)}_{i|cd}$ weakly converges to $\widetilde{\nu}_{i|cd}$ for $i=1,2$, respectively. Thus, (i) is proved.
We next prove (ii). For this, we compute the limits of $\varphi^{(r)}_{i}(0)$ for $i=1,2$ under the assumption in (ii). First consider the case that $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,t]$. In this case, $R^{(r)}_{s}(0-)$ has the same distribution as $T^{(r)}_{s}$ under $\mathbb{P}^{(r)}_{e}$, so, by Lemma 4.4 for $k=1$,
$\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)\;=\ell^{(r)}_{0})]$ $\;=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]+o(r).$
@@
$\quad=(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})(\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}])=o(r).$ (4.70)
Thus, if $b_{1}=0$, then it follows from (4.54) and (4.60) that
$\lambda_{2}\frac{\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
Since $\varphi^{(r)}_{1}(0)+\varphi^{(r)}_{2}(0)=1$ and $\mu c_{i}=\lambda_{i}$, this can be written as
$\frac{\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)=\frac{b_{2}}{c_{2}}(1-\varphi^{(r)}_{1}(0))+o(1).$
@@
On the other hand, if $b_{1}\not=0$, then, from (4.54) and (4.63),
$\frac{\lambda_{2}\mu b_{1}}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
Hence,

### v7

@@ 4.3 Expansion of the asymptotic BAR
From (4.3), we have
@@
$\mathbb{P}[L^{(r)}=0]$ $=r\frac{c_{1}\mu\sigma_{1}^{2}\big{(}\theta_{1}-\beta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})+2e^{\theta_{1}\ell_{0}}(b_{1}\mu\varphi^{(r)}_{1}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})}{2c_{1}\mu(e^{\theta_{1}\ell_{0}}-1)}+o(r).$ (4.54)
If $b_{1}=0$, then $\beta_{1}=b_{1}=0$, and therefore (4.54) yields
$\mathbb{P}[L^{(r)}=0]$ $=r\lim_{\theta_{1}\to 0}\frac{\theta_{1}(c_{1}\mu\sigma_{1}^{2}\psi^{(r)}_{1}(\theta_{1})+2e^{\theta_{1}\ell_{0}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})}{2c_{1}\mu(e^{\theta_{1}\ell_{0}}-1)}+o(r)$
$=\frac{r}{2c_{1}\mu\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),$
@@
which proves (4.53) for $b_{1}=0$. If $b_{1}\not=0$, then substituting $\theta_{1}=\beta_{1}$ into (4.54) yields
$\mathbb{P}[L^{(r)}=0]=\frac{re^{\beta_{1}\ell_{0}}}{c_{1}\mu(e^{\beta_{1}\ell_{0}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r),$
which is (4.53) for $b_{1}\not=0$. ∎
We also compute $\alpha^{(r)}_{e}\Delta^{(r)}_{1}$ to be used later.
Lemma 4.10.
@@
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\begin{cases}\frac{r}{2\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),&b_{1}=0,\\ \frac{r}{e^{\beta_{1}\ell_{0}}-1}\Big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\frac{1}{c_{1}\mu}\beta_{1}e^{\beta_{1}\ell_{0}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\Big{)}+o(r),&b\not=0.\end{cases}$ (4.55)
This lemma is immediate from substituting $\mathbb{P}^{(r)}_{e}[L^{(r)}=0]$ of Lemma 4.9 into (4.49). We are now ready to compute $\psi^{(r)}_{i}(\theta_{i})$ for $i~{}1,2$, which will be a key to prove Theorem 3.1.
Lemma 4.11.
@@
$\psi^{(r)}_{1}(\theta_{1})$ =$\begin{cases}\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}}\Big{(}\varphi^{(r)}_{1}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}}+o(1),&b_{1}=0,\\ \frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\Big{(}\varphi^{(r)}_{1}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}}+o(1),&b_{1}\not=0,\end{cases}$ (4.56)
$\psi^{(r)}_{2}(\theta_{2})$ $=\frac{\beta_{2}e^{\theta_{2}\ell_{0}}}{\beta_{2}-\theta_{2}}\Big{(}\varphi^{(r)}_{2}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}e^{\theta_{2}\ell_{0}}+o(1).$ (4.57)
Proof.
We first compute $\psi^{(r)}_{1}(\theta_{1})$. Assume that $b_{1}=0$. Substituting $\mathbb{P}^{(r)}_{e}[L^{(r)}=0]$ of Lemma 4.9 into (4.3) for $\beta_{1}=b_{1}=0$, we have
$-rc_{1}\mu\sigma_{1}^{2}\theta_{1}\psi^{(r)}(\theta_{1})$
$\quad=(1-e^{\theta_{1}\ell_{0}})\frac{r}{\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+r2e^{\theta_{1}\ell_{0}}\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r),$
which proves (4.56) for $b_{1}=0$. Next assume that $b_{1}\not=0$. Similarly to the case of $b_{1}=0$, it follows from Lemma 4.9 and (4.3) for $b_{1}\not=0$ that
$\frac{r}{2}c_{1}\mu\sigma_{1}^{2}(\beta_{1}-\theta_{1})\psi^{(r)}(\theta_{1})$
$\quad=\frac{re^{\beta_{1}\ell_{0}}(1-e^{\theta_{1}\ell_{0}})}{(e^{\beta_{1}\ell_{0}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+re^{\theta_{1}\ell_{0}}\big{(}b_{1}\mu\varphi^{(r)}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r)$
$\quad=\frac{rb_{1}\mu\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{0}}-1}\big{(}e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}})+r\frac{\beta_{1}e^{\beta_{1}\ell_{0}}(1-e^{\theta_{1}\ell_{0}})+\theta_{1}e^{\theta_{1}\ell_{0}}(e^{\beta_{1}\ell_{0}}-1)}{(e^{\beta_{1}\ell_{0}}-1)}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r).$
Hence, we have
$\psi^{(r)}_{1}(\theta_{1})$ $=\frac{\beta_{1}\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}-\theta_{1}e^{\theta_{1}\ell_{0}}+(\theta_{1}-\beta_{1})e^{\theta_{1}\ell_{0}}e^{\beta_{1}\ell_{0}}}{c_{1}\mu\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)(\beta_{1}-\theta_{1})}2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
$=\frac{\beta_{1}\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}$
$\quad+\frac{\beta_{1}(e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}})+(\beta_{1}-\theta_{1})e^{\theta_{1}\ell_{0}}-(\beta_{1}-\theta_{1})e^{\theta_{1}\ell_{0}}e^{\beta_{1}\ell_{0}}}{c_{1}\mu\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)(\beta_{1}-\theta_{1})}2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\Big{(}\varphi^{(r)}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2e^{\theta_{1}\ell_{0}}}{c_{1}\mu\sigma_{1}^{2}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
Thus, we have (4.56) for $b_{1}\not=0$. We next compute $\psi^{(r)}_{2}(\theta_{2})$. From (4.3), we have
$\psi^{(r)}_{2}(\theta_{2})=\frac{e^{\theta_{2}\ell_{0}}}{\beta_{2}-\theta_{2}}\Big{(}\beta_{2}\varphi^{(r)}_{2}(0)+(\beta_{2}-(\beta_{2}-\theta_{2}))\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}\Big{)}+\theta_{2}o(r),$
which proves (4.57). ∎
5 Proof of Theorem 3.1
From Lemma 4.11, we can see that, if
$\lim_{r\downarrow 0}{\mathcal{E}}^{(r)}_{1}=\lim_{r\downarrow 0}{\mathcal{E}}^{(r)}_{2}=0,$ (5.1)
then the limits of $\psi^{(r)}_{1}(\theta_{1})$ and $\psi^{(r)}_{1}(\theta_{2})$ can be obtained as long as the limits of $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{1}(0)$ exist. This suggests that (5.1) must be a key condition for Theorem 3.1. Taking this into account, we separate the proof of Theorem 3.1 in two steps by two propositions.
Proposition 5.1.
In Theorem 3.1, the assumption (3.f) can be replaced by (5.1).
Proposition 5.2.
Under the assumptions (3.a)–(3.e), (5.1) holds if
$\lim_{r\downarrow 0}({\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2})=0.$ (5.2)
Furthermore, (5.2) holds if (3.f) holds. Hence, (3.f) implies (5.1).
In the next two subsections, we prove these propositions. For their proofs, we will use the following lemmas, which are proved in Appendix A.5, respectively.
Lemma 5.1.
Under the assumptions (3.a)–(3.e), if
$\lim_{r\downarrow 0}\mathbb{P}[L^{(r)}(0)=\ell^{(r)}_{0}]=0,$ (5.3)
then
$\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0.$ (5.4)
Lemma 5.2.
Under the assumptions (3.a)–(3.e), if (5.4) holds, then
$\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{2}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0.$ (5.5)
Remark 5.1.
From the proof of Proposition 5.2, we can see that (5.4) and (5.5) imply (5.2), and therefore (5.3) is sufficient for (5.2) and (5.1). We think that (5.3) is likely true because $L^{(r)}(\cdot)$ would weakly converge to a reflected diffusion and the stationary distribution of the reflected diffusion has a density. However, the verification of (5.3) would need a process limit for $L^{(r)}(\cdot)$, so this verification is beyond the present framework.
5.1 Proof of Proposition 5.1
Throughout this proof, the part (i), (ii) and (iii) of Theorem 3.1 under the condition (5.1) instead of (3.f) are denoted by the part (i)’, (ii)’ and (iii)’, respectively.
We first prove the part (i)’. Since the condition in the part (i)’ is given by (3.6), Corollary 4.3 implies that
@@
By this formula, we can immediately see that Lemma 4.11 and (5.1) yield
@@
Since $\widetilde{\varphi}_{i|cd}$ is the moment generating function of $\widetilde{\nu}_{i|cd}$, (5.7) and (5.8) are equivalent to that $\nu^{(r)}_{i|cd}$ weakly converges to $\widetilde{\nu}_{i|cd}$ for $i=1,2$, respectively. Thus, the part (i)’ is proved.
We next prove the part (ii)’. For this, we compute the limits of $\varphi^{(r)}_{i}(0)$ for $i=1,2$ under the assumptions in the part (ii)’. First consider the case that $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,t]$. In this case, $R^{(r)}_{s}(0-)$ has the same distribution as $T^{(r)}_{s}$ under $\mathbb{P}^{(r)}_{e}$, so, by Lemma 4.4 for $k=1$,
$\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)\;=\ell^{(r)}_{0})]$ $\;=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]+o(r).$
@@
$\quad=(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})(\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}])=o(r).$ (5.9)
Compute $\alpha^{(r)}_{e}\Delta^{(r)}_{1}$ by Lemma 4.10 and (5.1), and recall (4.50), then
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\begin{cases}\frac{r}{2\ell_{0}}c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+o(r),&b_{1}=0,\\ \frac{r}{e^{\beta_{1}\ell_{0}}-1}b_{1}\mu\varphi^{(r)}_{1}(0)+o(r),&b\not=0,\\ \end{cases}$ (5.10)
$\alpha^{(r)}_{e}\Delta^{(r)}_{2}=r\mu b_{2}\varphi^{(r)}_{2}(0)+o(r).$ (5.11)
Then, for $b_{1}=0$, (5.1) implies that
$\lambda_{2}\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
Since $\varphi^{(r)}_{1}(0)+\varphi^{(r)}_{2}(0)=1$ and $\mu c_{i}=\lambda_{i}$ for $i=1,2$, this can be written as
$\frac{\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)=\frac{b_{2}}{c_{2}}(1-\varphi^{(r)}_{1}(0))+o(1).$ (5.12)
@@
Similarly, for $b_{1}\not=0$, from (5.10) and (5.11), we have
$\frac{\lambda_{2}b_{1}\mu}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$ (5.14)
Hence, $\mu c_{i}=\lambda_{i}$ for $i=1,2$ yields